# SimPSI: A Simple Strategy to Preserve Spectral Information in Time Series Data Augmentation

### v1

@@ Introduction
Time series data, whether univariate or multivariate, plays a crucial role in various domains such as medicine Lipton et al. (2016), physiology Jia et al. (2020), and sensory devices Yao et al. (2017). Unfortunately, it is limited to collecting data samples under consideration of different types, constraining the performance and capabilities of neural networks that learn from it. To address this issue, data augmentation Iwana and Uchida (2021); Um et al. (2017) is employed as a simple yet effective solution via artificially increasing the number of samples based on a slight variation or perturbation on the original samples.
Data augmentation techniques have been extensively studied for time series, incorporating methods such as Jittering, Scaling, Magnitude warping, Time warping, Permutation Um et al. (2017), Shifting Woo et al. (2022), and Dropout Yang and Hong (2022). These perturbations have been popular choices in the time domain. The data augmentation is also considered in the frequency domain via applying the Fourier transform to time series data. The spectrum is then randomly perturbed before being converted back into the time domain through the inverse Fourier transform. Notable techniques in this category include Frequency masking, Frequency mixing Chen et al. (2023), and Frequency adding Zhang et al. (2022).
@@
We have discovered that while the aforementioned data augmentation techniques show effectiveness in certain specific tasks Um et al. (2017), they do not generalize well to time series classification benchmarks. Our experimental evidence in Fig. 1 presents the ungeneralized effectiveness of data augmentation techniques according to the datasets, such as signal demodulation, human activity recognition, and sleep stage detection. [^1] Detailed information about the tasks and our experimental setup can be found in the Experiments section. Those techniques, though reliant on randomness, operate under the assumption that the core information within the data is preserved. However, the result suggests that perturbing the original time series data is heuristic and depends on the data domain, which leads to losing essential information necessary to solve the tasks.
@@
In this paper, we introduce a simple strategy for preserving spectral information during time series data augmentation, which we refer to as SimPSI. Our strategy involves mixing the original spectral data and its augmented form, weighted by a preservation map. After applying any time series data augmentation technique, SimPSI converts the original and augmented time series to the frequency domain. It then combines the original spectrum with the augmented version based on the weightage given by the preservation map, which indicates the importance score for each frequency component. The combined spectrum is subsequently transformed back to the time domain, resulting in the final output of our framework. The remaining efforts concentrate on defining a well-structured preservation map. We propose three types of preservation maps: magnitude spectrum, saliency map, and spectrum-preservative map. The first two types use the given data’s magnitude spectrum and saliency map Simonyan, Vedaldi, and Zisserman (2013) as the preservation map. For the spectrum-preservative map, we developed a preservation map generator that takes input spectrum data and returns the preservation map. This map is learned through a preservation contrastive loss function that influences differentiated model output scores based on the preservation quality. We also propose a training strategy for improved optimization. To demonstrate the efficacy of SimPSI, we apply it to various time series data augmentation techniques and compare performance across different benchmarks. We also create a simulation to assess whether the proposed method correctly identifies spectral regions to preserve during data augmentation. Our experimental results demonstrate that SimPSI significantly enhances the effectiveness of time series data augmentation techniques by preserving essential spectral information, thereby preventing unintentional loss of core spectral details.
@@
Various data augmentation techniques have been proposed for time series. One prevalent and intuitive strategy involves slightly altering the magnitude. For instance, Jittering Um et al. (2017) introduces additive white Gaussian noise, Scaling Um et al. (2017) multiplies by a random scalar value, Shifting Woo et al. (2022) adds a random scalar value, Magnitude warping Um et al. (2017) multiplies by a random polynomial curve, and Dropout Yang and Hong (2022) masks random time indices. An alternative approach involves modifying the time scale rather than the magnitude. Time warping Um et al. (2017), for instance, interpolates the time scale with a random polynomial curve, while Permutation Um et al. (2017) rearranges the time order. An additional method involves perturbing the spectrum. Techniques such as Frequency masking, Frequency mixing Chen et al. (2023), and Frequency adding Zhang et al. (2022) serve as simple strategies that appropriately perturb global dependencies in the time domain.
@@
Data augmentation inherently introduces perturbations into the original data. If not appropriately managed, these perturbations could lead to significant information loss or an introduction of unnecessary noise and ambiguity. To mitigate this, studies have focused on information preservation. In the vision domain, KeepAugment Gong et al. (2021) employs a saliency map Simonyan, Vedaldi, and Zisserman (2013) of each image to identify and preserve informative regions during augmentation. AugMix Hendrycks et al. (2020) generates a composite of various augmented views of the data and mixes it with the original data, weighted by a random scalar. This ensures the final image is not overly distanced from the original one. In natural language processing, SSMix Yoon, Kim, and Park (2021) and SMSMix Yoon et al. (2022) leverage saliency map to retain certain word sequences, ensuring that crucial information remains intact during the data augmentation process. For time series data, Input smoothing Liu et al. (2022) scales high-frequency entries in the frequency domain by a random scalar, thereby reducing the impact of data noise. However, its application is limited to noise reduction, and the degree of reduction is randomly determined.
@@
where $Norm$ is a min-max normalization so that values of the magnitude spectrum $P_{mag}\in\mathbb{R}^{L}$ are between 0 and 1. Preserving frequencies with large magnitudes makes the original and augmented data look alike, but the core information for solving the task might disappear. For instance, detecting abnormalities in the Electrocardiogram (ECG) signals relies on capturing the pattern of small high-frequency components Tragardh and Schlegel (2006), whereas the magnitude spectrum $P_{mag}$ eliminates the core frequencies for the classification during the data augmentation process just because those have a small magnitude.
@@
where $\theta_{G}$ and $\theta_{p}$ are the parameters of the preservation map generator $G$ and the classifier $\hat{p}$, respectively. Note that $\theta_{G}$ does not descend towards the gradient of $L_{cl}$. This prevents $G$ from learning undesirable local minima, such as returning the uniform scalar value across different frequencies or the same map across different samples, in which the preservation map is not adaptive to the input time series but acts as a uniform band-pass filter. Also, the preservation map generator $G$ is removed during inference, so $G$ updated by the classification loss might interrupt classifier training.
@@
We verified if the proposed method improves the performance of existing data augmentation techniques by capturing important spectral regions and preserving them. To do that, we devised a simulation where information is carried on a set of known frequencies. Inspired by the wireless communication domain Ryu and Choi (2023), we constructed a synthetic dataset by modulating a sequence of random bits into the corresponding frequencies of a signal, called frequency shift keying (FSK), and the task is demodulating it. We used 8 and 32 different frequencies for modulation (FSK8 and FSK32), and each dataset consists of 2,304 training signals, 288 validation signals, and 288 testing signals. We chose ResNet1D Ramjee et al. as a baseline network, which was given a 128-length modulated signal and returned a 32-length M-ary (M=8, 32) sequence. We used the Adam optimizer Kingma and Ba (2015) with the learning rate $10^{-3}$, and the networks were trained for 50 epochs. The training was performed on a single NVIDIA RTX A6000 GPU. Appendix A provides more details about our experimental setup.
@@
Scale-Shift-Jittering Woo et al. (2022) 94.6 $\pm$ 0.2 −0.2 92.8 $\pm$ 0.5 −1.2 80.7 $\pm$ 0.2 0
@@
Frequency masking Chen et al. (2023) 94.2 $\pm$ 0.3 −0.6 93.9 $\pm$ 1.7 −0.1 81.0 $\pm$ 0.4 +0.3
@@
Model 3-layer 2-layer 2-layer CNN LSTM Transformer Jittering Um et al. (2017) 95.1 92.2 94.2 + SimPSI ($P_{sp}$) 95.2 93.7 96.4
@@
In the HAR dataset Reyes-Ortiz et al. (2012), data is collected by the accelerometer and gyroscope of a smartphone mounted on a waist and sampled at 50 Hz, which aims to classify human activities. Following the data preprocessing in Eldele et al. (2021), an input time series has a length of 128 and nine channels. The dataset consists of 7,352 training samples and 2,947 test samples labeled with six classes. We chose a 3-layer CNN model for classification, which was used in Wang, Yan, and Oates (2017); Eldele et al. (2021); Zhang et al. (2022), and additionally included a 2-layer LSTM model and a 2-layer Transformer model for further verifications. We used the Adam optimizer with the learning rate $10^{-3}$, and the networks were trained for 100 epochs. We adhered to configurations in Eldele et al. (2021), and the training was performed on a single NVIDIA RTX A6000 GPU. Appendix B and C provide more details about our experimental setup.
@@
We used the SleepEDF dataset Goldberger et al. (2000) for classifying sleep stages from Electroencephalogram (EEG) signals sampled at 100 Hz. We followed the data preprocessing in Eldele et al. (2021), where the input has a length of 3,000 and a single channel. The dataset comprises 35,503 training samples and 6,805 test samples labeled with five classes. We chose a 3-layer CNN model for classification, also used in the human activity recognition experiments. We used the Adam optimizer with the learning rate $10^{-3}$, and the networks were trained for 40 epochs. We adhered to configurations in Eldele et al. (2021) for other details. The training was performed on a single NVIDIA RTX A6000 GPU.
@@
We used the Waveform dataset Moody (1983) for classifying rhythm types from ECG recordings of human subjects with atrial fibrillation. It was sampled at 250 Hz, and we followed the data preprocessing step as in Tonekaboni, Eytan, and Goldenberg (2021). Every input has a length of 2,500 and two channels. The dataset comprises 59,922 training samples and 16,645 test samples labeled with four classes. We chose a 1-dimensional strided CNN with six convolutional layers and a total down-sampling factor 16, proposed in Tonekaboni, Eytan, and Goldenberg (2021). We used the Adam optimizer with the learning rate $10^{-4}$, and the networks were trained for 8 epochs. We adhered to configurations in Tonekaboni, Eytan, and Goldenberg (2021) for other details. The training was performed on a single NVIDIA RTX A6000 GPU. Performance enhancement through SimPSI is described in Appendix D.
@@
Model $Accuracy$ $AUPRC$ Scale-Shift-Jittering 92.0 $\pm$ 1.9 62.6 $\pm$ 2.3 + SimPSI ($P_{sp}$) 95.2 $\pm$ 0.3 64.7 $\pm$ 2.0 + SimPSI ($P_{sp}$) w/o $L_{pc}$ 94.9 $\pm$ 0.1 63.7 $\pm$ 2.1 + SimPSI ($P_{sp}$) w/ joint training 94.5 $\pm$ 0.3 62.9 $\pm$ 1.9
@@
We ablated the proposed method from two perspectives, verifying the impact of the preservation contrastive loss and separate training strategy. We showed the performance of a 6-layer CNN model on the Waveform dataset while a composition of Scaling, Shifting, and Jittering Woo et al. (2022) was applied (Table 3). Removing the preservation contrastive loss resulted in a 0.3 decrease in accuracy and a 1.0 decrease in AUPRC. Applying joint training of the cross-entropy loss and the preservation contrastive loss made a 0.7 decrease in accuracy and a 1.8 decrease in AUPRC.

### v2

@@ Introduction
Time series data, whether univariate or multivariate, plays a crucial role in various domains such as medicine (Lipton et al. 2016), physiology (Jia et al. 2020), and sensory devices (Yao et al. 2017). Unfortunately, it is limited to collecting data samples under consideration of different types, constraining the performance and capabilities of neural networks that learn from it. To address this issue, data augmentation (Iwana and Uchida 2021; Um et al. 2017) is employed as a simple yet effective solution via artificially increasing the number of samples based on a slight variation or perturbation on the original samples.
Data augmentation techniques have been extensively studied for time series, incorporating methods such as Jittering, Scaling, Magnitude warping, Time warping, Permutation (Um et al. 2017), Shifting (Woo et al. 2022), and Dropout (Yang and Hong 2022). These perturbations have been popular choices in the time domain. The data augmentation is also considered in the frequency domain via applying the Fourier transform to time series data. The spectrum is then randomly perturbed before being converted back into the time domain through the inverse Fourier transform. Notable techniques in this category include Frequency masking, Frequency mixing (Chen et al. 2023), and Frequency adding (Zhang et al. 2022).
@@
We have discovered that while the aforementioned data augmentation techniques show effectiveness in certain specific tasks (Um et al. 2017), they do not generalize well to time series classification benchmarks. Our experimental evidence in Fig. 1 presents the ungeneralized effectiveness of data augmentation techniques according to the datasets, such as signal demodulation, human activity recognition, and sleep stage detection. [^1] Detailed information about the tasks and our experimental setup can be found in the Experiments section. Those techniques, though reliant on randomness, operate under the assumption that the core information within the data is preserved. However, the result suggests that perturbing the original time series data is heuristic and depends on the data domain, which leads to losing essential information necessary to solve the tasks.
@@
In this paper, we introduce a simple strategy for preserving spectral information during time series data augmentation, which we refer to as SimPSI. Our strategy involves mixing the original spectral data and its augmented form, weighted by a preservation map. After applying any time series data augmentation technique, SimPSI converts the original and augmented time series to the frequency domain. It then combines the original spectrum with the augmented version based on the weightage given by the preservation map, which indicates the importance score for each frequency component. The combined spectrum is subsequently transformed back to the time domain, resulting in the final output of our framework. The remaining efforts concentrate on defining a well-structured preservation map. We propose three types of preservation maps: magnitude spectrum, saliency map, and spectrum-preservative map. The first two types use the given data’s magnitude spectrum and saliency map (Simonyan, Vedaldi, and Zisserman 2013) as the preservation map. For the spectrum-preservative map, we developed a preservation map generator that takes input spectrum data and returns the preservation map. This map is learned through a preservation contrastive loss function that influences differentiated model output scores based on the preservation quality. We also propose a training strategy for improved optimization. To demonstrate the efficacy of SimPSI, we apply it to various time series data augmentation techniques and compare performance across different benchmarks. We also create a simulation to assess whether the proposed method correctly identifies spectral regions to preserve during data augmentation. Our experimental results demonstrate that SimPSI significantly enhances the effectiveness of time series data augmentation techniques by preserving essential spectral information, thereby preventing unintentional loss of core spectral details.
@@
Various data augmentation techniques have been proposed for time series. One prevalent and intuitive strategy involves slightly altering the magnitude. For instance, Jittering (Um et al. 2017) introduces additive white Gaussian noise, Scaling (Um et al. 2017) multiplies by a random scalar value, Shifting (Woo et al. 2022) adds a random scalar value, Magnitude warping (Um et al. 2017) multiplies by a random polynomial curve, and Dropout (Yang and Hong 2022) masks random time indices. An alternative approach involves modifying the time scale rather than the magnitude. Time warping (Um et al. 2017), for instance, interpolates the time scale with a random polynomial curve, while Permutation (Um et al. 2017) rearranges the time order. An additional method involves perturbing the spectrum. Techniques such as Frequency masking, Frequency mixing (Chen et al. 2023), and Frequency adding (Zhang et al. 2022) serve as simple strategies that appropriately perturb global dependencies in the time domain.
@@
Data augmentation inherently introduces perturbations into the original data. If not appropriately managed, these perturbations could lead to significant information loss or an introduction of unnecessary noise and ambiguity. To mitigate this, studies have focused on information preservation. In the vision domain, KeepAugment (Gong et al. 2021) employs a saliency map (Simonyan, Vedaldi, and Zisserman 2013) of each image to identify and preserve informative regions during augmentation. AugMix (Hendrycks et al. 2020) generates a composite of various augmented views of the data and mixes it with the original data, weighted by a random scalar. This ensures the final image is not overly distanced from the original one. In natural language processing, SSMix (Yoon, Kim, and Park 2021) and SMSMix (Yoon et al. 2022) leverage saliency map to retain certain word sequences, ensuring that crucial information remains intact during the data augmentation process. For time series data, Input smoothing (Liu et al. 2022) scales high-frequency entries in the frequency domain by a random scalar, thereby reducing the impact of data noise. However, its application is limited to noise reduction, and the degree of reduction is randomly determined.
@@
where $Norm$ is a min-max normalization so that values of the magnitude spectrum $P_{mag}\in\mathbb{R}^{L}$ are between 0 and 1. Preserving frequencies with large magnitudes makes the original and augmented data look alike, but the core information for solving the task might disappear. For instance, detecting abnormalities in the Electrocardiogram (ECG) signals relies on capturing the pattern of small high-frequency components (Tragardh and Schlegel 2006), whereas the magnitude spectrum $P_{mag}$ eliminates the core frequencies for the classification during the data augmentation process just because those have a small magnitude.
@@
where $\theta_{G}$ and $\theta_{p}$ are the parameters of the preservation map generator $G$ and the classifier $\hat{p}$, respectively. Note that $\theta_{G}$ does not descend towards the gradient of $\mathcal{L}_{cl}$. This prevents $G$ from learning undesirable local minima, such as returning the uniform scalar value across different frequencies or the same map across different samples, in which the preservation map is not adaptive to the input time series but acts as a uniform band-pass filter. Also, the preservation map generator $G$ is removed during inference, so $G$ updated by the classification loss might interrupt classifier training.
@@
We verified if the proposed method improves the performance of existing data augmentation techniques by capturing important spectral regions and preserving them. To do that, we devised a simulation where information is carried on a set of known frequencies. Inspired by the wireless communication domain (Ryu and Choi 2023), we constructed a synthetic dataset by modulating a sequence of random bits into the corresponding frequencies of a signal, called frequency shift keying (FSK), and the task is demodulating it. We used 8 and 32 different frequencies for modulation (FSK8 and FSK32), and each dataset consists of 2,304 training signals, 288 validation signals, and 288 testing signals. We chose ResNet1D (Ramjee et al. 2020) as a baseline network, which was given a 128-length modulated signal and returned a 32-length M-ary (M=8, 32) sequence. We used the Adam optimizer (Kingma and Ba 2015) with the learning rate $10^{-3}$, and the networks were trained for 50 epochs. The training was performed on a single NVIDIA RTX A6000 GPU. Appendix A provides more details about our experimental setup.
@@
Scale-Shift-Jittering (Woo et al. 2022) 94.6 $\pm$ 0.2 −0.2 92.8 $\pm$ 0.5 −1.2 80.7 $\pm$ 0.2 0
@@
Frequency masking (Chen et al. 2023) 94.2 $\pm$ 0.3 −0.6 93.9 $\pm$ 1.7 −0.1 81.0 $\pm$ 0.4 +0.3
@@
Model 3-layer 2-layer 2-layer CNN LSTM Transformer Jittering (Um et al. 2017) 95.1 92.2 94.2 + SimPSI ($P_{sp}$) 95.2 93.7 96.4
@@
In the HAR dataset (Reyes-Ortiz et al. 2012), data is collected by the accelerometer and gyroscope of a smartphone mounted on a waist and sampled at 50 Hz, which aims to classify human activities. Following the data preprocessing in (Eldele et al. 2021), an input time series has a length of 128 and nine channels. The dataset consists of 7,352 training samples and 2,947 test samples labeled with six classes. We chose a 3-layer CNN model for classification, which was used in (Wang, Yan, and Oates 2017; Eldele et al. 2021; Zhang et al. 2022), and additionally included a 2-layer LSTM model and a 2-layer Transformer model for further verifications. We used the Adam optimizer with the learning rate $10^{-3}$, and the networks were trained for 100 epochs. We adhered to configurations in (Eldele et al. 2021), and the training was performed on a single NVIDIA RTX A6000 GPU. Appendix B and C provide more details about our experimental setup.
@@
We used the SleepEDF dataset (Goldberger et al. 2000) for classifying sleep stages from Electroencephalogram (EEG) signals sampled at 100 Hz. We followed the data preprocessing in (Eldele et al. 2021), where the input has a length of 3,000 and a single channel. The dataset comprises 35,503 training samples and 6,805 test samples labeled with five classes. We chose a 3-layer CNN model for classification, also used in the human activity recognition experiments. We used the Adam optimizer with the learning rate $10^{-3}$, and the networks were trained for 40 epochs. We adhered to configurations in (Eldele et al. 2021) for other details. The training was performed on a single NVIDIA RTX A6000 GPU.
@@
We used the Waveform dataset (Moody 1983) for classifying rhythm types from ECG recordings of human subjects with atrial fibrillation. It was sampled at 250 Hz, and we followed the data preprocessing step as in (Tonekaboni, Eytan, and Goldenberg 2021). Every input has a length of 2,500 and two channels. The dataset comprises 59,922 training samples and 16,645 test samples labeled with four classes. We chose a 1-dimensional strided CNN with six convolutional layers and a total down-sampling factor 16, proposed in (Tonekaboni, Eytan, and Goldenberg 2021). We used the Adam optimizer with the learning rate $10^{-4}$, and the networks were trained for 8 epochs. We adhered to configurations in (Tonekaboni, Eytan, and Goldenberg 2021) for other details. The training was performed on a single NVIDIA RTX A6000 GPU. Performance enhancement through SimPSI is described in Appendix D.
Model $Accuracy$ $AUPRC$ Scale-Shift-Jittering 92.0 $\pm$ 1.9 62.6 $\pm$ 2.3 + SimPSI ($P_{sp}$) 95.2 $\pm$ 0.3 64.7 $\pm$ 2.0 + SimPSI ($P_{sp}$) w/o $\mathcal{L}_{pc}$ 94.9 $\pm$ 0.1 63.7 $\pm$ 2.1 + SimPSI ($P_{sp}$) w/ joint training 94.5 $\pm$ 0.3 62.9 $\pm$ 1.9
@@
We ablated the proposed method from two perspectives, verifying the impact of the preservation contrastive loss and separate training strategy. We showed the performance of a 6-layer CNN model on the Waveform dataset while a composition of Scaling, Shifting, and Jittering (Woo et al. 2022) was applied (Table 3). Removing the preservation contrastive loss resulted in a 0.3 decrease in accuracy and a 1.0 decrease in AUPRC. Applying joint training of the cross-entropy loss and the preservation contrastive loss made a 0.7 decrease in accuracy and a 1.8 decrease in AUPRC.